Instrument Log 

AERI X           C1                 CF Lamont, OK Optical Trailer

Date

Operator

Comments

2/29/2000

CAM

There was no graph displayed on the AERIX ingest computer. All graphs and status boxes were displayed on the FEP computer but they were not current. The last time displayed on the graphs and in the status window was 0715 GMT. Pressed the reset button on the AERIX electronics. The ingest computer began updating and the graph appeared. Selected the shut down option on the FEP computer and rebooted it at 1459 GMT. The system then came up normally and all graphs were present and current. RTS status O.K. at 1507 GMT.

2/28/00

KT

At 2000 GMT reset the instrument switch and the ingest computer started scrolling and producing a graph. The status boxes never changed their appearance. They were either red, blue, or green. Troubleshooting still in progress.

2/23/2000

CAM

John Van Allen contacted Site Ops and informed that he was unable to access AERIX computers. T/S found that the FEP computer was at the desktop display with no graphs visible. The ingest computer was locked up with no graph displayed. Performed a graceful shut down and restart of the FEP computer and power cycled the ingest computer at 1630 GMT. Reset the AERIX electronics. System then booted up normally and all graphs returned. Filled the AERIX with LN2 and RTS status O.K. at 1655 GMT.

2/22/00

KT

The ingest computer was locked up. At 2115 GMT pressed the reset and attempts failed. At 2117 GMT rebooted system. After several attempts, still no change. Troubleshooting still in progress.

2/16/2000

CAM

John Van Allen contacted Site Ops and informed that he was unable to access AERIX computers. T/S found that the FEP computer was at the desktop display with no graphs visible. The ingest computer was locked up with no graph displayed. Neither computer could be shut down gracefully. Powered both computers off and reset the AERIX electronics. Powered computers on. ScanDisk found and corrected several disk errors on the FEP computer. System then booted up normally and all graphs returned. Filled the AERIX with LN2 and RTS status O.K. at 1700 GMT.

2/14/2000

KT

There are no graphs of any kind that are displayed, as well as status boxes that are not displayed. System is at desktop.

2/4/2000

Mentor

Instrument back on line.

2/3/2000

Mentor

John Olson on site for AERIX maintenance.

8/12/99

MR

Talked to John Van Allen and he had me confirm that the white light source was burnt out. He said that this was something they would have to replaced since it involves calibration and alignment of the light afterwards. He said to shut the unit down but leave the dry purge air connected and on. He did not give any indication as to when they would repair the unit. He also requested that the SyJet disk drive be removed and sent back to them for repairs since the unit will be down and this would be a good time to look at the drive. Turned off all power and removed the Syjet drive.

8/10/99

MR

T/S indicated that the constant white light source inside the unit has gone out. Awaiting response back from John Van Allen.

8/9/99

MR

Shut down the instrument due to hardware problems. Mentor says he will arrange for a tech to arrive on site to perform maintenance and repairs.

8/2/99

MR

Rebooted the system due to no graphs being displayed. This did not work. Called Conner Flynn and left voice message to call back and the status of the AERIX.

7/16/99

MR

Removed the plastic covered cardboard that was covering the hole.

7/6/99

MR

The DOS machine is stopping with the following error: Not ready reading drive E: Rebooted the both computers at 1430 GMT. After rebooted the DOS machine took one sample and stopped running again with the same error. Performed a ScanDisk check on the E: drive. The check found some corrupted files in the file 070699 and corrected them. It then suggested running a surface scan on the disk. Performed the surface scan. This still did not repair the problem. Called Fred Best and left a voice message for him to call.

4/15/99

MK

The system computer does not have any data displayed. The FEP computer is displaying all of the proper graphs but no data is present. The ENCODER SCENE CONFIRM is in condition red, with a value of -999. Shutdown and rebooted both the FEP and AERI X system computer at 1420 GMT. Also reset the AERI X electronics. No changed has yet to be noticed in either computers.

4/12/99

MK      

The FEP monitor was only displaying the desktop icons. The AERI X system monitor was locked up and was displaying an error message unable to read drive E. The SYQUEST drive has ejected the cartridge. Rebooted both the FEP and AERI X system computers at 1515 GMT. Inserted the cartridge back in to the SYQUEST drive. RTS at 1545 GMT.

4/8/99

CAM

John Van Allen sent a procedure for changing the AERIX Syjet 1.5GB disk. Changed the disk out according to procedure at 1900 GMT. The procedure called for executing sqshell.exe and formatting the new disk from the menu. Upon executing sqshell.exe the ingest computer rebooted automatically and resumed normal operation without opening a shell and allowing the user to format the disk. Checked the Syjet drive (E:) and found that all necessary files and directories were on the new disk, so the disk appears to have formatted. RTS at 2000 GMT. Will send old disk to John Van Allen.

3/23/99

MR

Rebooted computer due to no graphs being displayed. Tried to notify the mentor but there was no answer. Will try to contact mentor later today.

3/16/99

MW

The AERIX monitor still shows no signal and there are no graphs or status boxes on the FEP display. Pressed the reset button several times but this did not correct the problem. Rebooted the system computer at 1723 GMT. Still no signal. Will continue to let the system run to see if it recovers.

3/15/99

MW

The CF lost power due to a snow storm on 3/14/99 at 0030 GMT. The AERIX has been down since that time. The hatch was found in the open position and snow had fallen into the enclosure and partially obscured the mirror assembly. Cleared the snow inside and outside of the enclosure. Cleaned the mirror with distilled water and isopropyl alcohol. Powered the system back up at 2030 GMT after AC power was restored. Could not get a signal on the AERIX display and as a result the graphs and status boxes would not come up on the FEP display. Pressed the AERIX reset button several times to no avail. Will let the system run as-is overnight to see if it recovers.

2/19/99

SS

No graphs displayed on the FEP monitor. Reset the AERIX electronics at 1600 GMT. Status ok. RTS 1620 GMT.

1/21/99

MW

The AERIX hatch was not closing all the way. Oiled the hinge areas with WD40 and tested operation 2000-2015 GMT. Status OK.

1/13/99

MW

The AERIX scan LED is flashing but the AERIX ingest computer displays no signal. The FEP has stopped collecting data. The date displayed on the FEP status box window is 10101 (not a typo). The time displayed on the FEP status box window is UTC 04:07:53. Pressing the reset button on the AERIX has no effect. Rebooting the system has no effect. The instrument mentor is out of the country until 1/16/1999. Will leave system as is for now to see if it recovers on its own.

1/5/99

MW

No graphs displayed on the FEP monitor. No signal displayed on the DOS machine monitor. Rebooted the FEP at 1652 GMT. Pressed the AERIX reset button at 1654 GMT and the signal returned. The FEP came up with graph boxes displayed. Had to press the AERIX reset button again at 1742 GMT because the signal disappeared. The signal returned after pressing the reset button several times. Returned to service with the signal and graphs displayed at 1800 GMT.

12/28/98

MW

The AERIX FEP and DOS machine displays indicated no signal from the AERIX. There were no display boxes on the FEP and there was no signal on the DOS machine. Shut down the FEP and DOS machine at 1638 GMT. Power cycled the AERIX. Restarted the FEP and then the DOS machine. The DOS machine began displaying the signal but the signal appears to be very noisy. No graphs appeared on the FEP. Contacted the instrument mentor and he suggested rebooting the system again. Shut down the FEP and DOS machine at 1845 GMT. Power cycled the AERIX. Restarted the FEP and then the DOS machine. No signal present on the DOS machine and no graphs displayed on the FEP after reboot. Pressed the AERIX reset button at 1916 GMT and the signal showed up on the DOS machine display. The FEP began displaying graph display boxes but the system locked up with a large grey box displayed in front of the graph display boxes. Rebooted the FEP at 1939 GMT. The FEP came back on-line at 1941 GMT with graph boxes displayed. However, the graphs do not appear to be updating. At approximately 2020 GMT, the signal disappeared from the DOS machine display. The DOS machine rebooted itself around 2028 GMT. No signal was displayed. Pressed the AERIX reset button once again at 2033 GMT. The signal reappeared. The FEP is still not updating. Contacted the instrument mentor and informed him of the problem. He indicated that the attenuated signal might mean that something is out of alignment and not enough data is being sent to the FEP to generate calibrated graph displays. The mentor suggested letting the system run as is until he can dispatch some personnel to correct the problem.

12/22/98

MW

The AERIX did not appear to be scanning according to the display. No graphs were displayed on the FEP. The Syquest drive was full and the system locked up when the DOS machine tried to write to the Syquest drive. Placed another disk in the Syquest drive and the FEP graph windows came back up after rebooting both the FEP and DOS machine. However, according to the DOS machine display, the AERIX did not appear to be scanning and no data was displayed in the FEP graph windows. The AERIX scan LED was flashing normally as if it was scanning. Rebooted the system several times, also attempted using the AERIX reset button several times to no avail. Called the instrument mentor and left a message. Decided to press the reset button one last time before leaving the optical trailer at 2100 GMT. Suddenly the AERIX DOS display began scanning normally. Shortly thereafter, data began showing up on the FEP graph windows. Not sure if the problem is with the AERIX, the DOS computer HW/SW, or the communications link between the two machines. RTS at 2110 GMT.

12/21/98

KT

Both of the AERI X computers showed no signs of instrument working. The graph computer was showing the windows icon screen and the data scrolling computer had an end code. At 1836 GMT rebooted both computers. No change. At 1838 GMT rebooted systems once again and still no change. At 1840 GMT rebooted system again and also the AERI X instrument and there was still no change. Attempted to get in touch with Mr. Van Allen which had no success. Instrument was left in current state until further notice.

11/9/98

SS

The mentor is not receiving data from the instrument. Found the FEP computer locked up at 00:16:41 on 11/09/98 and the system computer continually scrolls Error: Unable to resolve host 'dorothy.sgp.arm.gov'. Rebooted both computers at 1739 GMT. Both computers came up and appear to be functioning properly. RTS at 1745 GMT.

11/5/98

CAM

Instrument mentor advised that the AERIX DOS computer was locked up and needed reset. Power cycled the AERIX DOS computer at 1750 GMT. Reset the electronics at 1757 GMT. RTS status O.K. at 1800 GMT.

10/16/98

TG

The AERIX computer was not visible on the network. The application used previously to set TCP/IP had improperly configured the config.sys file. Corrected config.sys file and loaded the correct protocol.ini file. RTS at 1645 GMT. The system and FEP computers were then able to communicate with each other and the network. However the AERIX instrument was not communicating with the computer. Reset AERIX electronics at 1650 GMT and instrument is now operational. Returned the hatch to auto mode and RTS status O.K. at 1652 GMT.

10/15/98

MW

The instrument mentor repaired one of the circuit boards for the AERIX and requested that it be installed. Filled the AERIX with LN2 and inserted the repaired circuit board at 1830 GMT. Powered the system on but the AERIX computer could not find the network. Notified site computer operations and they contacted the instrument mentor for further instructions. The AERIX will be down until the network settings can be corrected.

10/9/98

ZW

The AERI-X is being worked upon today by the Mentor.

10/09/98

TG

SDS requested that the aerix-fep computer address be changed from 198.124.97.41 to 198.124.98.72. Powered up computer and Changed IP Address to 198.124.98.72. Changed Router address to 198.124.98.1, Moved Fiber connections from 97 Hub to 98 Hub, Rebooted computer and pinged address 198.124.97.105. John VanAllen was able to ping aerix-fep computer from his computer. Status O.K. Shutdown computer.

9/24//98

KT

System is still powered down. Hatch is still closed until further notice.

9/21/98

KT

System is still powered down. Hatch mode was taken out of auto and placed into manual and the hatch opening was closed until further notice.

9/14/98

SS

Instrument mentor called and request that the instrument be powered down until he can have it repaired. Turned the power of to the instrument and shut down both computers.

9/9/98

SS

Instrument mentor requested that the instrument be reset after it has had time to cool down. Reset the instrument numerous times with out being able to get the instrument to start a scan. RTS at 1815 GMT.

9/9/98

SS

Instrument mentor called and asked that the instrument be turned back on and the Dewar filled. Turned the power on to the FEP and system computer and instrument at 1535 and filled the Dewar at 1540 GMT. RTS at 1600 GMT.

8/25/98

SS

Morning rounds revealed that the AERIX system computer was idle with the message "Driver not found press any key to continue. Pressed space bar and the computer completed boot up. But locked up with the same message again when checked at 1448 GMT. Left a message for Mr. John Van Allen explaining problem. Mr. John Van Allen returned my call at 1511 and determined that the system will need to be turned off until the instrument specialists returns from a field trip. Shut down the computers and secured the instrument power.

8/20/98

MR

Daily PM found no display on either computer monitors. Rebooted both computers.

8/11/98

MK

The FEP was not displaying any graphs or status boxes and the AERIX computer display had no signal. Reset electronics at 1535 GMT. System computer now displays graph and appears to be properly running.

8/5/98

ZW

During morning rounds found AERI-X with no status boxes displayed, and no graphs were displayed at 1411 GMT. At 1450 GMT Site Ops hit the reset switch. The status boxes, and the graphs returned to normal at 1510 GMT. Status ok. RTS 1510 GMT.

7/30/98

MW

The FEP was not displaying any graphs or status boxes and the AERIX computer display had no signal. The AERIX electronics scan LED was flashing normally. Reset the AERIX electronics several times with no effect. Rebooted both computers at 1425 GMT but this didn't help either. Powered the entire system down at 1440 GMT and powered it back up at 1443 GMT. No effect. Reset the AERIX electronics four more times at 1540 GMT and scanning resumed on the fourth reset. Rebooted the FEP at 1540 GMT. The FEP came up with graphs and status boxes at 1555 GMT. RTS.

7/1/98

MW

The air compressor which feeds air to the AERIX has malfunctioned and is not running. The compressor motor tripped its motor circuit breaker. When the motor circuit breaker was reset, the motor tripped the AC power circuit breaker. After the AC power breaker was reset, the motor would not activate. T/S in progress. Informed the AERIX mentor of the problem.

6/23/98

CAM

During daily PM checks the FEP monitor was at the desktop icon view and their was no scan graph on the AERIX monitor. Rebooted both computers and reset electronics at 1500 GMT. System RTS with good data at 1515 GMT.

6/16/98

SS

Found the AERIX FEP computer not displaying status and the system computer not displaying a graph. Reset electronics at 1553 GMT. System computer now displays graph and appears to be properly running program. RTS at 1600 GMT.

6/2/98

MK

Due to a power outage at 2030 GMT on 6/1/98 the AERIX system had failed to restart. At 1409 GMT on 6/2/98, rebooted the AERIX computer, the FEP computer and the AERIX electronics box. At 1459 GMT all status boxes had returned to normal. RTS at 1500 GMT.

5/29/98

CAM

After removing the insulation panels to clean the AERI01 scene mirror it was noted that the AERIX mirror also needed cleaned. Stopped the AERIX program using ctrl-k and q at 1534 GMT. Placed the scene mirror in the sky position at 1535 GMT. Used electronics vacuum cleaner to remove bugs and grass from within the enclosure surrounding the scene mirror. Cleaned dust build up from the scene mirror using distilled water and isopropyl alcohol. Rebooted AERIX computer at 1539 GMT. RTS at 1549 GMT.

5/26/98

MK

Due to a power failure on 5/25/98 at 0425 GMT, the FEP and AERI computers are both down. Shutdown and restarted both computers at 1452 GMT. status OK. RTS at 1530 GMT.

5/18/1998

SS

Instrument mentor indicated that the instrument had stopped and could not be restarted remotely. Completed a graceful shut down of both computers and rebooted at 1645 GMT. Reset electronics after computers came up. RTS at 1715 GMT.

5/15/98

SS

Due to power failure restarted instrument at 1525 GMT.

5/7/98

SS

Found the system not running. Reset the FEP and system computer and system electronics at 1326 GMT. RTS at 1345 GMT.

5/5/98

MW

The AERIX was not scanning. There was no signal on the monitor and no graph on the FEP. The 'Encoder Scene Confirm' status box was red. Reset the AERIX electronics at 1509 GMT. Scanning resumed immediately after the reset. The FEP returned to normal status by 1520 GMT. RTS.

5/1/98

MW

The AERIX FEP had no status boxes displayed since yesterday. Shut down the FEP and rebooted at 1548 GMT. RTS at 1550 GMT with all graphs and status boxes displayed.

4/27/98

MW

Water was slowly dripping on top of the AERIX electronics enclosure. It appeared to be leaking where the external housing meets the optical trailer. Cleaned up all excess water and placed a tin tub underneath the drip. By 1630 GMT, the dripping had almost completely subsided.

4/27/98

MW

The AERIX computer and FEP were locked up this morning. The AERIX computer screen displayed the message 'Terminate batch job?' followed by a 'Y/N?' prompt. The FEP displayed the message 'The system has halted.' Power cycled the AERIX electronics, terminated the batch job on the AERIX computer and rebooted the AERIX computer and FEP at 1510 GMT. The system came back online with the AERIX computer scanning and graphs displayed on the FEP at 1630 GMT. RTS.

4/10/98

MR

The AERIX had stopped scrolling and the program was hung up in a software loop. Rebooted both systems by pressing the CTRL\ALT\DEL keys. System came up and started to scan at 1541. Still no graphs or status boxes displayed on the FEP.

4/6/98

MW

The FEP was not displaying graphs or status boxes. Power cycled the FEP and AERI computer at 1720 GMT. System came back online with graphs and status boxes at 1730 GMT.

3/20/98

SS

Hatch manually opened at 1449 GMT, mode selector placed in auto position. RTS at 1449 GMT.

3/19/98

SS

The hatch was manually closed at 1540 due to snow and blowing snow. The moisture sensor was not being activated by the snow.

3/12/98

SS

Found the system computer scrolling but not displaying a graph or updating. The FEP computer was not running the AERI program. 1439 Reset the electronics. 1507 The system computer appears to have good graphs, the FEP computer has a graph and the statuses are all green with the exception of HBB Temp Stability value -1e+10, HBB Max Temp Diff value 0.534576, and HBB NEN value nan. RTS at 1515 GMT.

3/11/98

SS

Found the system and FEP computers not running the AERI program. 1424 Re-set the electronics. 1427 The system computer began running the AERI program. 1435 The FEP computer began running the AERI program with all green except the HBB Max Temp (yellow) and HBB NEN (blue). 1454 both computer quit taking data. 1504 Reset electronics. 1558 the system computer is running the AERI program but is only intermittently showing good graphs. The FEP computer is not showing a graph but is updating with status. RTS at 1600 GMT.

3/10/98

MW

The AERIX was not scanning and no waveform was displayed on the screen. Reset the AERIX electronics at 1935 GMT. Status OK after reset.

3/10/98

SS

Reset the electronics at 1447 GMT now the system computer is running with a graph and data. Rebooted both the FEP and System computers at 1501 GMT. At 1507 GMT the FEP computer came up with data and status screens no graph and mostly blue statuses. At 1509 GMT the system computer came up with graph and started scrolling. At 1523 GMT the statuses went green on all but the HBB Max Temp Diff. RTS at 1523 GMT.

3/9/98

SS

Found FEP monitor not displaying the status or graph and the system monitor continuously scrolling. With error "unable to find F:AERI.DO. Reset electronics no help. Reset both FEP and System computer no help. The insulation for the AERI was blown in and un-stabilized the room temperature and may be affecting the AERIX. Called John Van Allen unable to reach him will attempt reset after room temperature has stabilized tomorrow morning.

3/5/98

SS

During daily checks found internal insulation panels in need of slight adjustment. Made adjustments and secured panels with tape. RTS 1515 at GMT.

3/3/98

SS

Found the FEP computer in TELEX mode. Was unable to restart status screen. Rebooted computer at 1550 GMT. Computer came up and displayed a good graph and all status indicators are green. Found the system computer not displaying a graph. Reset the system electronics at 1538. The computer came up and had a good graph. RTS at 1555 GMT.

2/25/98

CM/MW

The instrument mentor requested that Site Ops check the AERI01 view port for obstructions. After removing the insulation panels and cleaning the AERI01 scene mirror it was noted that the AERIX scene mirror had bird droppings on it. Stopped the AERIX program at 1705 GMT. Placed the mirror in the sky position and cleaned the mirror using distilled water and isopropyl alcohol. Replaced the insulation panels and rebooted AERIX computer. RTS at 1715 GMT.

2/6/98

MW

The instrument mentor called site ops and said that the AERIX was not operating properly. Noticed no signal on the AERIX display at 2125 GMT. Reset the AERIX electronics at 2127 GMT. The signal returned. RTS.

1/16/98

MW

The instrument mentor asked site ops to change the temperature set point level of the AERI black body temperature controller from 960 to 985. Changed the set point from 960 to 985 at 1850 GMT.

1/12/98

MW

The instrument mentor asked site ops to change the temperature set point level of the AERI black body temperature controller from 985 to 960 to compensate for colder winter conditions. Changed the set point from 985 to 960 at 1629 GMT.

1/5/98

MW

There was no graph displayed on the AERIX computer screen. The FEP graphs and tool bars were not displayed either. Reset the AERIX electronics at 1617 GMT. The graph returned on the AERIX display. Soft booted the FEP at 1621 GMT but the graphs and tool bars did not come up. Hard booted the FEP at 1630 GMT. The normal FEP display came up at 1635 GMT. RTS.

12/30/97

MW/CM

Noticed the AERIX mirror was slightly dirty when the insulation panels were removed to clean the AERI01 mirror. Placed the AERIX mirror in the 'sky' position at 1558 GMT. Cleaned the mirror with isopropyl alcohol and water at 1600 GMT. Rebooted the AERIX computer at 1603 GMT. Replaced the insulation panels at 1604 GMT. RTS at 1605 GMT.

12/11/97

CAM

The system monitor spectrum continues to be intermittent. The graph appears to be O.K. on blackbody scans, but disappears or becomes noisy on sky scans. Also several status boxes are blue or yellow. Taped some of the insulation panels to better seal against outside winds. The indoor thermostat cool setting was raised from 70F to 74F and the heat setting was raised from 68F to 72F at 1620 GMT. This was done at mentors request to see if it helps with some of the AERIX temperature problems. Checked system status again at 1830 GMT and the changes appear to have helped. The system monitor spectrum graph was consistent and the blue status boxes had cleared.

12/10/97

CAM

While performing maintenance on another system in the optical trailer, it was discovered that AERIX was locked up. The last good data sample received on the FEP was at 1516 GMT. There were several blue status boxes on the FEP and there was no spectrum displayed on the system monitor. System status had been checked previously and filled with LN2 at 1515 GMT and was operational at that time. Rebooted both computers at 1605 GMT. When operation resumed there was still no spectrum displayed on the system monitor and the FEP data sample time and date indicated 0. Performed a complete system shut down and restart at 1618 GMT. The FEP computer found and corrected several allocation file errors. Returned to check status again at 1702 GMT and everything appeared normal. However, one more check at 1830 GMT found that the spectrum display on the system monitor was intermittent. At times the spectrum looked normal, but would then disappear. At other times the spectrum appeared as noise the entire length of the graph. Notified mentor of the situation and left in present state. Mentor is looking into possible causes and will get back with us.

12/01/97

CAM

John Olson contacted Site Ops and suggested that we try pressing the ‘reset’ button on the J2 card in the instrument electronics box. Pressed the reset button several times and the instrument finally resumed communications with the system computer at 1730 GMT. The FEP displays then updated at 1737 GMT and system returned to normal operations.

11/26/97

CAM

There was no graphical data present on the Mean Downwelling Radiance graph and several of the status boxes were blue. Rebooted system at 1456 GMT. System came back up at 1501 GMT with no change in status. Power cycled the electronics and rebooted system again at 1508 GMT. This time the FEP UTC time came up 00:00:00 and the date showed 0 as well. Also several status boxes were red with values of 0 and the remaining status boxes were blue indicating no data available. Further attempts to reboot the system proved unsuccessful with the status remaining unchanged and time and date showing 0. It appeared as though the PC was unable to communicate with the instrument. Was unable to contact mentor due to the holiday, but spoke with Frank Murcray at Denver University who advised he would attempt to locate someone for support. System left in present state for now.

11/20/97

CAM

There was no data displayed on the Mean Downwelling Radiance. Also several of the status boxes were blue, indicating there was no data available. Rebooted system at 1510 GMT. RTS status O.K. at 1522 GMT.

11/12/97

CAM

The Responsivity, Detector Cooled, and HBB Nen status boxes were all red with values of 0.0101288, 1.48077, and 85.021 respectively. Also the 'Mean Down Welling Radiance' graph on the FEP was nearly all red. The system was not filled with LN2 on 11/11 due to the holiday. Filled with LN2 at 1558 GMT. At 1603 GMT the red cleared from the graph and the Detector Cooled status box returned to green. HBB Nen status box was then blue and HBB Temp Stability turned red but returned to green at 1619 GMT when the detector temperature stabilized after filling with LN2.

10/31/97

CAM

During the daily PM checks the status box display was not visible on the AERIX FEP computer. Shut down the AERIX FEP and system computers at 1455 GMT. Rebooted both computers and RTS at 1501 GMT. Status boxes then came up O.K. The beam splitter base temperature status box is yellow with a value of 299.887.

10/2/97

MW

Cleaned the AERIX mirror today while the insulation panels were removed to clean the AERI01 mirror. Interrupted data acquisition and placed the mirror in the "sky" position at 1519 GMT. Cleaned the mirror with distilled H2O and isopropyl alcohol at 1520 GMT. Rebooted the AERIX computer at 1521 GMT. RTS at 1550 GMT.

9/25/97

MW

The AERIX computer had no graph displayed and the FEP status boxes did not change to a green condition after the LN2 fill this morning. Reset the AERIX electronics at 1418 GMT. The AERIX computer still did not display a graph and the FEP displayed no response. Rebooted the AERIX computer and FEP at 1432 GMT. Still no graph displayed on the AERIX computer. Reset electronics again at 1436 GMT and the graph finally appeared on the AERIX computer monitor. The FEP returned to normal with all status boxes green at 1450 GMT.

9/22/97

MW

The AERIX "FEPX SurfaceLayerAirTemp," "FEPX LW_HBB_NEN," and "FEPX calibrationHBBtemp" graphs are displaying disjointed/scrambled looking data lines. No corrective maintenance performed at this time.

9/15/97

MW

The AERIX FEP and computer monitors did not have the graphs displayed. The AERIX did not appear to be scanning. Rebooted the system at 1445 GMT. The graphs did not come up on the FEP and the scanning graph did not appear on the AERIX computer monitor. Depressed the AERIX scanning electronics reset button at 1502 GMT. This brought up the scanning graph on the computer monitor but the FEP displays did not start. Rebooted the FEP at 1505 GMT. The data graphs popped up at 1513 GMT. RTS.

9/5/97

MW/CM

There was no problem with the AERIX. The mirror was cleaned while the insulation panels were off for maintenance on the AERI01 mirror. Exited the AERIX program at 1841 GMT and placed the mirror in the 'sky' position. Cleaned the mirror with distilled water and isopropyl alcohol 1842-1843 GMT. Rebooted the FEP and AERIX computer at 1844 GMT. RTS at 1910 GMT with all status boxes green.

9/3/97

MDW

Due to a Kay Electric transformer going down in Billings on 9/2, the site lost all power at around 2030 GMT. The extended power outage caused the optical trailer UPS to drain its batteries and all instruments lost power. Site power was restored at 0142 GMT on 9/3 but the AERIX did not come back up correctly and no graphical data was displayed. Rebooted ingest computer at 1426 GMT. Rebooted FEP computer at 1427 GMT. Still no data was displayed. Power cycled the mirror stepping motor at 1445 GMT and rebooted the ingest and FEP computers again. System RTS status O.K. at 1450 GMT.

9/3/97

MW

The interferometer enclosure RH status box was yellow as of 9/2/97. The air flow meter which feeds the AERIX was closed. Adjusted the air flow to the AERIX to 20 PSI for two minutes, then adjusted the airflow to 5 PSI. The status box returned to a green condition by 1515 GMT.

9/2/97

MW

The interferometer enclosure RH status box is yellow with a value of 15.6333. This is 0.6333 higher than allowed for a green condition. No corrective maintenance performed at this time.

8/22/97

MW/MR

Site technicians noticed that the AERIX mirror looked dirty when they were cleaning the AERI01 mirror. Moved the AERIX mirror to the 'sky' position at 1529 GMT. Cleaned the mirror at 1530 GMT. Rebooted the AERIX computer and FEP at 1534 GMT. RTS at 1555 GMT.

8/18/97

MR

Received a shipment of LN2. Filled AERIX with LN2 and restarted system at 1910 GMT. John Van Allen called at 2030 GMT and indicated that the mirror controller did not appear to be working. A check of the instrument found that all components were powered on, however John advised that the controller needs to be powered on before the ingest computer is powered on because the controller receives initialization commands from the PC. Reset the ingest computer at 2050 GMT and RTS status O.K.

8/15/97

MW

There is no LN2 in stock to fill the AERIX dewar so an instrument shutdown was performed. Powered the preamp, laser, compu-motor, and AERIX computer down at 1514 GMT. Left FEP powered up. Also shut both monitors off.

7/28/97

MW

The 'Responsivity', 'HBB Temp Stability', and 'Detector Cooled' status boxes were red. Added LN2 to dewar at 1401 GMT. RTS with all status boxes green at 1423 GMT

7/28/97

CM/CW

AC power to the optical trailer had to be shut off at 1845 GMT for construction. Performed graceful shut down at 1843 GMT. AC power restored at 1905 GMT and system RTS status O.K. at 1910 GMT.

7/17/97

CM/MW

During daily PM it was discovered that the hatch was inoperative. The hatch remained open during a heavy rainstorm and could not be closed manually. Covered the AERIX mirror with plastic at 1950 GMT until further T/S and repairs can be performed when storm passes.

7/17/97

MW/CM

T/S of hatch problem indicated voltage all the way up to the motor (12Vdc). R/R the DC motor at 1815 GMT. the original motor smelled as if it had burned up. While testing the operation of the new motor, it seemed that the limit switch which limits the downward motion of the hatch is not working as it should. If the limit switch is manually depressed as far as possible, the hatch will stop immediately. However, the oval wheel assembly does not push the limit switch down far enough to reach this stopping point. The motor continues to drive the hatch for a short period after the hatch has closed but it does not appear to be straining too much. Over a long period of time, this could possibly account for the original motor burning up. The limit switch which limits the upward motion of the hatch does not have to be depressed near as far as the downward switch does in order to engage - thus indicating a probable faulty downward motion limit switch. Do not have spare limit switch to replace at this time. The new motor allows the hatch to operate in auto mode for now. RTS at 1845 GMT.

7/17/97

CM/MW

Due to a failed hatch motor, the mirror got rained on momentarily before it could be covered. The mirror had water spots on it and needed to be cleaned. Exited the AERIX program at 1845 GMT and placed the mirror in the sky position. Cleaned the mirror using distilled water and isopropyl alcohol. Restarted system and RTS status O.K. at 1910 GMT.

7/15/97

CM/MW

The AERI01 mirror was cleaned today and at that time a decision was made to clean the AERIX mirror as well. Exited the AERIX program and placed the mirror in the "sky" position at 1927 GMT. Cleaned mirror using distilled water and isopropyl alcohol and vacuumed the vicinity using an electronics vacuum. RTS status O.K. at 1943 GMT.

7/1/97

MW

The status box window was not displayed on the FEP screen. Rebooted FEP at 1415 GMT. RTS at 1420 GMT.

6/30/97

MW

The AERIX had several status boxes which were blue and one status box was red. Blue status boxes: HBB Temp, ABB Temp, HBB NEN, HBB Max Temp Diff, ABB Max Temp Diff, HBB Temp Stability, Responsivity. Red Status Box: Detector Cooled. Added LN2 to dewar at 1640 GMT. All status boxes returned to green condition by 1703 GMT.

6/27/97

CAM

While cleaning the AERI01 mirror it was noticed that AERIX mirror also needed cleaned and there was significant debris (grass and bird droppings) inside the cavity surrounding the mirror. Halted AERIX program and cleaned the mirror using distilled water and isopropyl alcohol. Used an electronics vacuum cleaner to clean dirt and debris from around and inside the cavity. AERIX RTS status O.K. at 2020 GMT. Prior to cleaning the mirror was noted to be mostly covered with light dust.

6/16/97

MK

On daily rounds it was found that a red flag was displayed on the monitor Encoder Scene Confirm. This problem has reference to the mirror position. Power cycled the computer at 1535 GMT nothing changed power cycled electronics at 1617 also nothing changed. Observed mirror position and found no obstruction around the mirror. Status not ok. RTS 1632 GMT.

5/23/97

MR

System is down until further notice with hardware problems.

5/12/97

ZW

The graphs were not displayed on monitor. Power cycled the AERIX program at 1415 GMT to no avail. Called Raphael Dedecker who advised to try power cycling the electronics as well as the program. Power cycled electronics and program at 1548 GMT to no avail. Tried power cycling electronics and program several times with no positive result. Powered off the AERIX at 1630 GMT Mentor is aware of the problem and somebody from the University of Denver is coming to fix the problem.

5/08/97

MW

Opened hatch and placed in auto at 1424 GMT.

5/08/97

MW

The DRD11A Rain Detector sensor for the AERIX was damaged, possibly by birds. *R/R rain detector from 1530 GMT to 1535 GMT. New detector has a protective shield. RTS at 1540 GMT. * Old - PO7415 * New - S11101

5/07/97

MS

Closed AERI hatch at 2115 GMT due to expected rain overnight.

5/01/97

MW

The HBB temp stability status box was red. Added LN2 at 1412 GMT. All boxes green as of 1425 GMT.

4/30/97

ZW

Daily PM. The Monitor was showing three red boxes. Filled AERIX with LN2 at 1458 GMT. The boxes returned to green at 1530 GMT. Status ok. RTS 1530 GMT.

4/29/97

MW

Three red status boxes and one yellow status box. Added LN2 at 1355 GMT. Status boxes green as of 1415 GMT.

4/28/97

ZW

Mentor repaired the AERIX. The AERIX is operating normally as of 1300 GMT.

4/24/97

MW

The AERIX hatch was not responding to the control switches. Measured the output of the 12V power supply. The voltmeter indicated 3Vdc. Unhooked the load from the power supply and measured 12Vdc. Removed the cover of the hatch motor assembly at 1800 GMT. Wasps had built up nests around the motor power lines and around the oval wheels which trip the limit switches. Removed all debris/larvae from around the motor/drive assembly at 1810 GMT. Hooked the load back up to the power supply and attempted to open the hatch at 1820 GMT. The hatch would not open. Lifted the hatch a few inches by hand to clear the limit switch and the motor started to rotate slightly but stopped. Measured 8V across the motor. Unplugged the AC from the switch box and R/R the motor at 1830 GMT. Plugged the AC back in to the switch box and tested the hatch operation at 1845 GMT. The motor was straining to open the hatch but it could not clear the limit switches without manually helping it by hand. After helping the hatch clear the limit switch, it would open or close all the way. Sprayed WD-40 around the rusty hatch hinges at 1850 GMT. After lubricating the hinge, the hatch opened and closed normally without straining the motor. Placed the hatch in auto mode and RTS at 1900 GMT.

3/26/97

MLS

The AERI-X is still inoperative per Mentor.

3/17/97

CAM

John Van Allen advised there is a hardware problem with AERIX computer and system will be down indefinitely.

3/14/97

MS

During daily PM, the AERI-X’s computer was found locked up. Upon advising the Mentor, I was advised that the AERI-X would be out of operation for a little while today as he was doing some work on it. The AERI-X was again checked at 1900 GMT. The system was still down at that time.

2/25/97

CAM

Assisted mentor and received training with cleaning the optics mirror at 2200 GMT. AERI X RTS at 2230 GMT.

2/24/97

KT

At 1607 GMT the Detector cooled status box was highlighted in RED. At the time there was no LN2 to fill up the AERI X with. At 1725 GMT after filling the LN2 bottle up I proceeded to fill up the AERI X which in turn returned the Detector cooled status box back to condition green. Status is o.k. at this time. RTS at 1730 GMT.

2/11/97

CAM

During daily PM the Beam splitter Base Temp was condition yellow and HBB Max Temp Diff was condition red at 1600 GMT. Returned at 1645 GMT to perform LN2 fill and conditions had then returned to green.

2/10/97

MR

Filled LN2 and placed the hatch back in auto mode at 1508 GMT. Graph started to show data at about 1520 GMT. RTS at 1520 GMT.

2/03/97

CAM

John Olson is on site to work on instrument.

1/22/97

CAM

John Van Allen advised to power off instrument and ingest computer. Task completed at 1730 GMT. FEP computer left running for remote maintenance.

1/21/97

DKB

Daily check LN2 fill 1508 GMT. No data display or system status on FEP. Sys computer scrolls but does not seem to be collecting data. Informed mentor.

1/16/97

CAM

Ingest computer locked up. "Inet not Found" displayed. FEP data not updating. Hard boot ingest computer and soft boot FEP at 1520 GMT. Ingest computer would not load Windows 95. Reboot again at 1537 GMT. No graph displayed on ingest computer and FEP status boxes are all red or blue. Status uncertain at this time.

1/13/97

MR/KT

Mentor advised system was locked up and needed rebooted. Reboot and RTS at 1548 GMT.

1/09/97

GRS

No data displayed on either monitor. Mentor advised to power cycle instrument and reboot computers. Reboot at 1519 GMT. Status unknown with FEP at desktop icon view and no graph displayed on ingest computer. Mentor is looking into the problem.

12/24/96

CAM

No data displayed on either monitor. Rebooted both computers at 1625 GMT with no change. FEP at desktop icon view and no graph visible on system monitor. Left message for Dedecker and RTS as is.

12/23/96

CAM

PC clock was 1 hour slow. Reset clock and restarted PC at 1446 GMT.